OpenStudy (anonymous):

Calculate the equation of the line between the points (-6 -1) and (2 -1)

OpenStudy (marissalovescats):

What are you trying to do? Make a line equation from these 2 points?

OpenStudy (marissalovescats):

Well 1st you are going to use these 2 points to get the slope.

OpenStudy (anonymous):

No, all I want is the equation.

OpenStudy (marissalovescats):

I know. Sorry got kicked off. From there use the slop and your points to put into the point slope formula. y-y1=m(x-x1)

OpenStudy (marissalovescats):

And then put that into y=mx+b.

OpenStudy (anonymous):

Since the y value is the same at both points, we know this is a horizontal line. The equation of the line would seem to be\[y=-1\]